# What Is That Talk About? A Video-to-Text Summarization Dataset for Scientific Presentations

### v3

@@ 5 Experiments
We benchmark our dataset using three settings: zero-shot learning, QLoRA fine-tuning Dettmers et al. (2024), and full-parameter fine-tuning. For zero-shot, we test closed-source multimodal models, including GPT-o1 Achiam et al. (2023), Gemini 2.0 Team et al. (2023), Claude 3.5 Sonnet Anthropic (2024), as well as open-source video LMMs such as Video-LLaMA Zhang et al. (2023), Video-ChatGPT Maaz et al. (2024), Video-LLaVA Lin et al. (2024a), LLaMA-VID Li et al. (2025), LLaVA-NeXT-Interleave Li et al. (2024b), and mPLUG-Owl3 Ye et al. (2024). These open-source video LMMs process videos by extracting multimodal features, such as visual and/or audio components, using cross-modal attention mechanisms to align and integrate information across modalities.
We also assess LLaMA-3.1 and Qwen2-Audio to examine if text- or audio-based models can accomplish the summarization task without taking video information into account. For LLaMA-3.1, we explore two variants: in LLaMA-3.1transcript, we extract audio from video files using moviepy and transcribe it with OpenAI’s Whisper-1 to generate text input for the model. In LLaMA-3.1OCR, we apply EasyOCR to extract on-screen text from video frames and use the OCR-generated text as input for summarization. Similarly, for Qwen2-Audio, we use moviepy to convert video files into audio and treat the audio as input. Exact model versions are provided in Appendix D. Based on our benchmarking results, we select the best-performing model as the backbone for the plan-based strategy and evaluate its performance. Prompts for the above models are provided in Appendix J (Figures 1114).
@@
To ensure a fair comparison, all models, including baselines, plan-based models, and ablation models, are evaluated under identical hyperparameter settings unless explicitly stated otherwise. All models are tested using identical prompt instructions. Detailed hyper-parameter configurations are provided in Appendix E.
@@
We report a set of evaluation metrics to measure informativeness, alignment, and factual consistency in summaries. For informativeness, we use ROUGE Lin (2004), SacreBLEU Post (2018), METEOR Banerjee and Lavie (2005), BERTScore Zhang et al. (2020), and CIDEr-D Vedantam et al. (2015). Specifically, we provide the F1 scores for Rouge-1 (R1), Rouge-2 (R2), and Rouge-LSum (RLSUM). Alignment to the input video is evaluated with VideoScore He et al. (2024), and factual consistency with FactVC Liu and Wan (2023). Detailed descriptions of these metrics are given in Appendix F.
@@ 6 Results and Analysis
Table 3 illustrates the performance differences between closed-source and open-source models. In the zero-shot setting, closed-source models generally outperform their open-source counterparts. Among open-source models, mPLUG-Owl3 stands out, particularly in semantic alignment (BERTScore) and video-text consistency (VideoScore). Fine-tuning on in-domain data yields noticeable improvements for open-source models with both QLoRA and full-parameter fine-tuning. QLoRA shows overall lower performance than full parameter fine-tuning.
LLaMA-3.1transcript, LLaMA-3.1OCR, and Qwen2-Audio perform similarly on our dataset. While both text- and audio-based models achieve competitive results, video-based LMMs demonstrate overall superior performance, with mPLUG-Owl3 achieving SOTA results across most metrics. This result further underlines the importance of video for our summarization task.
@@
Plan-mPlug-Owl3 is the plan-based approach built on mPLUG-Owl3, outperforming all open-source baselines in both zero-shot and fine-tuned settings. For zero-shot inference, the Plan-mPlug-Owl3 variant, which fine-tunes only the Plan Generation (PG) module, surpasses other models in summary quality, factual consistency, and semantic alignment. With full-parameter fine-tuning, Plan-mPlug-Owl3 achieves the highest overall scores across models, showing improvements in factual accuracy (+3.47 in FactVC) and quality (+0.34 in RLsum) compared to mPLUG-Owl3. However, all models (including the plan-based method) exhibit hallucinations (FactVC) and alignment (VideoScore) issues, and there are still significant differences (p-value of the paired t-test is less than 0.05) between the human performance in this task, with reference abstracts scoring 88.54 on FactVC and 4.62 on VideoScore.
Impact of Plan Generation Strategy
@@
We analyze the plan generation ablation strategy by comparing it with simpler baselines: Lead-3Q, Tail-3Q, and Random-3Q. In these ablation baselines, plans are generated by selecting the first three, last three, or three randomly chosen summary sentences, respectively. Each selected sentence serves as a target for generating a question, with its preceding sentences providing the context. For instance, in the Lead-3Q setting, the first sentence is used as the target (without any preceding context), prompting the first question in the plan, while subsequent sentences incorporate earlier ones as context. Additionally, we compare the case where QUD is not considered. That is, we directly let GPT-o1 generate all planning questions at once only based on the reference summary (NoQUD).
@@
Table 4: Performance comparison of different plan generation strategies under full fine-tuning settings. Textual content at the start of the summary is more helpful for generating plans.
Table 4 underlines the performance differences across different plan generation ablation strategies. NoQUD is also a plan-based approach. It has lower data processing overhead than our original method and performs better than the end-to-end method. However, it still falls short to some extent compared to our approach. The Lead-3Q strategy performs better overall compared to Tail-3Q and Random-3Q, indicating that initial sentences offer stronger contextual continuity for generating plan questions. Nonetheless, these heuristic strategies fail to match the performance of the original planning method.
@@
We assess how the quality of the plan questions affects model performance. We applied GPT-o1 as a question generator in a zero-shot setting in our previous experiments. For comparative analysis, we additionally incorporate Llama-3.1 and a state-of-the-art question generation algorithm (RAST) from Gou et al. (2023) to generate the plan questions. In addition, we apply a Random Replacement (RR) method, where questions generated by GPT-o1 are randomly replaced with irrelevant ones. The number of replaced questions per summary ranges from one to the entire set. We also introduce full random replacement (FRR), where questions generated by GPT-o1 are all replaced with randomly irrelevant questions. [^3] The prompt for generating irrelevant questions is given in Appendix Figure 15.
@@
Figure 5: Noise in plan generation impacts summarization performance. FRR is a shorthand for Full Random Replacement and RR for Random Replacement. RAST is a SOTA question generation method.
@@
Figure 5 reveals that the quality of plan questions does influence the summarization performance: using GPT-o1 to generate questions outperforms the rest. The FRR method performs worst, as irrelevant questions disrupt the alignment between the plan and summary content. We also find that the plan-based method exhibits a certain degree of robustness, as it performs reasonably well even when the plans contain some degree of noise (RR vs. FRR). These findings emphasize the importance of question relevance and quality in structuring the output summaries. In Appendix G, we further explore the effect of video content on our summarization task, varying the length of the video given as input to the model. We also perform experiments with different textual contexts for generating plan questions, and with controlled generation. Additionally, we present an error analysis of model output in Appendix H, which highlights the gap between model-generated summaries and human-written references.
@@ 7 Human Evaluation
We conduct a human evaluation on 50 randomly selected instances from the VISTA test set. Annotators include master’s and doctoral students in computer science or computational linguistics with advanced English proficiency. They receive compensation per our university’s standard rate and are blind to the source of each summary to ensure impartial assessment. We compare Plan-mPlug-Owl3, mPLUG-Ow13, LLAVA-NeXT-Interleave, and GPT-o1 against human reference summaries/abstracts. Three independent annotators are asked to review the source video and evaluate corresponding model summaries (and the human upper bound) on a 1–5 Likert scale for Faithfulness, Relevance, Informativeness, Conciseness, and Coherence (higher scores indicate better quality). They are also asked to provide an overall ranking. In total, participants rated 750 samples ($50\times 5\times 3$). Appendix K contains the full annotation instructions.
@@
Figure 6 presents the performance of each model, along with the proportion of instances where models are rated best or worst. Fleiss’ Kappa scores for Faithfulness ($\kappa=0.767$), Relevance ($\kappa=0.842$), Informativeness ($\kappa=0.721$), Conciseness, and Coherence ($\kappa=0.813$) indicate a substantial level of agreement, with an average agreement score of $\kappa=0.787$. Overall, human-written summaries outperform all neural summarization models in quality, as they are perceived as substantially more faithful, coherent, concise, and informative. Human-written summaries are 81.7% more likely to be rated as best compared to model-generated summaries.
Among the four neural models, GPT-o1 performs worst, being rated as worst 63.2% of the time. LLAVA-NeXT-Interleave follows suit, with a 17.8% chance of receiving the worst ranking. The plan-based model, Plan-mPLUG-Owl3, outperforms mPLUG-Owl3 and demonstrates superior performance across all metrics. Additionally, it stands out among neural summarization systems for its higher likelihood of generating high-quality summaries. Paired t-tests show that human answers are considered significantly better than all neural models in all metrics ($p<0.05$), revealing a clear gap between automatic systems and human performance on the VISTA dataset. The plan-based method is significantly better ($p<0.05$) than other neural models in faithfulness, coherence, and informativeness, although it falls short of human performance. We also evaluate all samples of the test set with an LMM-as-Judge and obtain results that are broadly consistent with human evaluation. We describe the details of this study in Appendix I.
@@ 8 Conclusion
This paper introduces VISTA, a dataset for summarizing scientific video presentations into concise textual summaries. Comprehensive evaluations across multiple large models demonstrate that the summarization task is challenging, relying on the interplay of multiple modalities (video, text, and audio). We further introduce a plan-based approach, which yields improvements in summary quality and factual accuracy. Beyond dataset creation, our work also confirms that current leading large models still exhibit a noticeable gap compared to human performance.
@@
All the summary and video data used in this study are open source. While our sources are generally of high quality and exhibit a broad range of diversity, we have not investigated inherent biases in the data. Moreover, as these data represent only a small fraction of real-world data, our findings may not extend to all video-to-text summarization scenarios. In addition, our dataset is restricted to English, which limits its generalizability to other languages.
@@
In our task, we consider the paper abstract as the summary of the corresponding video. This hypothesis has been supported by our two-stage quality control process, which ensures a strong alignment. However, we acknowledge that there may be nuanced differences between the abstract and a textual summary derived solely from the video. That said, authors often present the abstract as a summary of the video, as it conveys the key contributions, objectives, and findings of the research, which are typically central to the content discussed.
@@
We use several state-of-the-art large models in our experiments and select the best-performing model, mPLUG-Owl3, to demonstrate the effectiveness of the planning strategy. These large models may carry biases introduced during pretraining. We have not assessed the extent of these biases, as they lie beyond the scope of this study. Furthermore, we have not tested the plan-based approach on all model variants presented in our experiments (e.g., text-based large models and audio-based large models). Our work does not aim to prove that the plan-based method is effective in all models of different modalities, but rather to demonstrate that the plan-based method can improve the performance of video-based models on our dataset. Moreover, plan-based methods can take many different forms, and our work does not aim to identify the optimal planning approach for our dataset. Future work could examine how the plan-based method performs across a wider range of models and modalities.
Modality
In our experiments, we explore the performance of individual modalities on the downstream summarization task (e.g., text-to-text model and audio-to-text model). However, we do not conduct an in-depth analysis of how different modality combinations impact the final summarization results. For instance, combining video transcripts with their visual content (transcript $+$ video-to-text) or with their audio (transcript $+$ audio-to-text) could yield different outcomes. Moreover, most video LMMs do not incorporate audio components; we also do not investigate how the integration of different modality components within video LMMs affects summarization results. The exploration of such modality combinations and their influence on summarization is left for future work.
Data Contamination and Prompt Selection
It is worth noting that we have not found evidence in the original papers describing the open-source models we use to suggest that the contents of the VISTA dataset are included in their pretraining stage. However, for closed-source models, such verification is not possible due to the lack of transparency in their pretraining datasets. Additionally, for the sake of consistency and fairness, we utilize the same prompts throughout our experiments, chosen primarily based on human judgment. However, since the number of possible prompts is limitless, other prompts could yield different outcomes. These factors represent potential directions for future research.
@@
LMM-as-Judge Evaluation
Although the LMM-based judge paradigm (GPT-o1) enables large-scale and relatively consistent evaluations, it may inherit biases from its pretraining data, and its black-box nature makes the rating process difficult to interpret. Data contamination also remains a concern if GPT-o1 is trained on overlapping data. We validate GPT-o1’s ratings with human evaluations on a small subset of samples, but this may not fully capture the model’s reliability across diverse topics, domains, or summary styles. Therefore, results should be interpreted with caution and supplemented by human judgment where possible.

### v4

@@ 5 Experiments
We benchmark our dataset using three learning settings: Zero-shot learning, QLoRA fine-tuning Dettmers et al. (2024), and full-parameter fine-tuning. For zero-shot learning, we test closed-source multimodal models, including GPT-o1 Achiam et al. (2023), Gemini 2.0 Team et al. (2023), Claude 3.5 Sonnet Anthropic (2024), as well as open-source video LMMs such as Video-LLaMA Zhang et al. (2023), Video-ChatGPT Maaz et al. (2024), Video-LLaVA Lin et al. (2024a), LLaMA-VID Li et al. (2024c), LLaVA-NeXT-Interleave Li et al. (2025), and mPLUG-Owl3 Ye et al. (2025). These open-source video LMMs process videos by extracting multimodal features, such as visual and/or audio components, using cross-modal attention mechanisms to align and integrate information across modalities.
We also assess LLaMA-3.1 Touvron et al. (2023) and Qwen2-Audio Chu et al. (2024) to examine if text- or audio-based models can accomplish the summarization task without taking video information into account. For LLaMA-3.1, we explore two variants: In LLaMA-3.1transcript, we extract audio from video files using moviepy and transcribe it with OpenAI’s Whisper-1 to generate text input for the model. In LLaMA-3.1OCR, we apply EasyOCR to extract on-screen text from video frames and use the OCR-generated text as input for summarization. Similarly, for Qwen2-Audio, we use moviepy to convert video files into audio and treat the audio as input. Exact model versions are provided in Appendix D. Based on our benchmarking results, we select the best-performing model as the backbone for the plan-based strategy and evaluate its performance. Prompts for the above models are offered in Appendix M (Figures 1114).
@@
To ensure a fair comparison, all models, including baselines, plan-based models, and ablation models, are evaluated under identical hyperparameter settings unless explicitly stated otherwise. All models are tested using identical prompt instructions. Detailed hyperparameter configurations are presented in Appendix E.
@@
We report a set of evaluation metrics to measure informativeness, alignment, and factual consistency in summaries. For informativeness, we utilize ROUGE Lin (2004), SacreBLEU Post (2018), METEOR Banerjee and Lavie (2005), BERTScore Zhang et al. (2020), and CIDEr-D Vedantam et al. (2015). Specifically, we provide the F1 scores for Rouge-1 (R1), Rouge-2 (R2), and Rouge-LSum (RLSUM). Alignment to the input video is evaluated with VideoScore He et al. (2024), and factual consistency with FactVC Liu and Wan (2023). Detailed descriptions of these metrics are given in Appendix F.
@@ 6 Results and Analysis
Table 3 compares model performance across three learning settings: Zero-shot, QLoRA fine-tuning, and full-parameter fine-tuning. Overall, fine-tuning on in-domain data yields substantial performance gains across all evaluation metrics. Full fine-tuning consistently outperforms QLoRA. While closed-source models such as GPT-o1 and Gemini typically lead in zero-shot performance, open-source models like mPLUG-Owl3 and Plan-mPlug-Owl3 achieve competitive or even superior results when fine-tuned, especially in semantic alignment (BERTScore) and video-text consistency (VideoScore).
We observe that video-based LMMs consistently outperform text-based and audio-based models. While models such as LLaMA-3.1transcript, LLaMA-3.1OCR, and Qwen2-Audio yield comparable results, they lag behind video-grounded models in overall performance. In particular, mPLUG-Owl3 achieves SOTA results across most metrics, highlighting the crucial role of visual information in enhancing summarization quality.
@@
Plan-mPlug-Owl3 is the plan-based approach built on mPLUG-Owl3, outperforming all open-source baselines in both zero-shot and fine-tuned settings. For zero-shot inference, the Plan-mPlug-Owl3 variant, which fine-tunes only the Plan Generation (PG) module, surpasses other models in summary quality, factual consistency, and semantic alignment. With full-parameter fine-tuning, Plan-mPlug-Owl3 achieves the highest overall scores across models, showing improvements in factual accuracy (+3.47 in FactVC) and quality (+0.34 in RLsum) compared to mPLUG-Owl3. However, all models (including the plan-based method) exhibit hallucinations (FactVC) and alignment (VideoScore) issues, and there are still significant differences (p-value of the paired t-test is less than 0.05) between the human performance in this task, with reference summaries scoring 88.54 on FactVC and 4.62 on VideoScore.
Impact of Modality Interplay
To explore the impact of different modality combinations on our multimodal tasks, we conduct an experiment using Video-LLaMA Zhang et al. (2023). Seven modality combinations are considered, including unimodal inputs (video, audio, transcript) and their pairwise or joint combinations. For each configuration, only the corresponding modality modules are updated while the remaining ones are kept frozen. The summarized results are shown in Table 4.
Modality Zero-shot Learning QLoRA Fine-tuning Full Fine-tuning
R2 RLsum VideoScore FactVC R2 RLsum VideoScore FactVC R2 RLsum VideoScore FactVC
Video only 2.68 20.34 1.55 28.93 8.83 27.51 2.65 50.66 10.78 30.02 2.91 60.87
Audio only 2.14 19.72 1.41 26.84 7.52 26.34 2.48 45.79 9.23 27.93 2.73 58.02
Transcript only 2.02 18.01 1.34 25.53 6.91 24.33 2.39 44.87 8.44 25.81 2.35 54.11
Video + Audio 3.19 21.24 1.63 32.25 9.44 28.33 2.77 52.05 11.86 31.68 3.04 64.21
Video + Transcript 1.87 18.94 1.39 27.76 7.35 24.82 2.51 48.63 9.01 27.19 2.65 58.91
Audio + Transcript 1.64 18.55 1.35 27.48 7.23 24.73 2.38 47.15 8.57 25.82 2.54 55.39
Video + Audio + Transcript 1.92 19.13 1.47 28.60 7.37 25.29 2.52 50.72 9.22 27.21 2.61 59.30
Table 4: Performance comparison of different modality combinations.
The results consistently show that video is the strongest standalone modality, likely due to its rich spatial-temporal information. Audio offers complementary prosodic and timing cues, but lacks semantic visual grounding. The transcript, while semantically rich, often introduces long, noisy, and unstructured inputs, particularly from ASR systems, that can overwhelm the model’s attention and interfere with alignment. These findings suggest that current video-based LMMs face challenges in effectively aligning and fusing token-heavy, noisy textual inputs with corresponding visual or audio information.
Impact of Plan Generation Ablations
@@
We analyze the plan generation ablation by comparing it with simpler baselines: Lead-3Q, Tail-3Q, and Random-3Q. In these ablation baselines, plans are generated by selecting the first three, last three, or three randomly chosen summary sentences, respectively. Each selected sentence serves as a target for generating a question, with its preceding sentences providing the context. For instance, in the Lead-3Q setting, the first sentence is used as the target (without any preceding context), prompting the first question in the plan, while subsequent sentences incorporate earlier ones as context. Additionally, we compare the case where QUD is not considered. That is, we directly let GPT-o1 generate all plan questions at once based on the reference summary (NoQUD).
@@
Table 5: Performance comparison of different plan generation ablations under full fine-tuning settings.
Table 5 underlines the performance differences across different plan generation ablations. For NoQUD, it underperforms compared to the QUD-based approach. The Lead-3Q strategy performs better overall compared to Tail-3Q and Random-3Q, indicating that initial sentences offer stronger contextual continuity for generating plan questions.
@@
We assess how the quality of the plan questions affects model performance. We apply GPT-o1 as a question generator in a zero-shot setting in our previous experiments. For comparative analysis, we additionally incorporate Llama-3.1 and a state-of-the-art question generation algorithm (RAST) from Gou et al. (2023) to generate the plan questions. In addition, we apply a Random Replacement (RR) method, where questions generated by GPT-o1 are randomly replaced with irrelevant ones. The number of replaced questions per summary ranges from one to the entire set. We also introduce full random replacement (FRR), where questions generated by GPT-o1 are all replaced with random irrelevant questions. [^2] The prompt for generating irrelevant questions is given in Appendix Figure 15.
@@
Figure 5: Noise in plan generation impacts summarization performance. FRR is a shorthand for Full Random Replacement, and RR for Random Replacement. RAST is a SOTA question generation method.
@@
Figure 5 reveals that the quality of plan questions does influence the summarization performance: Using GPT-o1 to generate questions outperforms the rest. The FRR method performs the worst, as irrelevant questions disrupt the alignment between the plan and summary content. We also find that the plan-based method exhibits a certain degree of robustness, as it performs reasonably well even when the plans contain some degree of noise (RR vs. FRR). These findings emphasize the importance of question relevance and quality in structuring the output summaries.
Planning Beyond Vision
While our primary objective is to evaluate the planning framework in the context of video-to-text summarization, it is valuable to assess its applicability to unimodal, non-visual models. To this end, we conduct supplementary experiments applying the planning method to three models that do not utilize video inputs: (1) LLaMA-3.1transcript (ASR-based textual input), (2) LLaMA-3.1OCR (OCR-based textual input), and (3) Qwen2-Audio (audio-based input). For each model, we compare baseline performance (i.e., without planning) against the planning counterpart. As summarized in Table 6, planning consistently improves performance across all settings and evaluation metrics. A paired t-test confirms that these improvements are statistically significant ($p<0.05$).
These findings demonstrate that the planning method does not function solely as a domain-specific enhancement but rather as a generalizable scaffold that supports better discourse structure, even in the absence of visual input. We hypothesize that, for text- and audio-based models, planning mitigates the lack of spatial-temporal signals by providing discourse-level anchors, such as intent-driven prompts (e.g., “What problem is being addressed?”), that guide the model’s summarization trajectory.
Model Setting R2 RLsum VideoScore FactVC
LLaMA-3.1transcript Zero-shot Learning 4.22 $\rightarrow$ 4.56 21.39 $\rightarrow$ 22.01 1.53 $\rightarrow$ 1.75 34.32 $\rightarrow$ 40.78
QLoRA Fine-tuning 11.38 $\rightarrow$ 11.62 30.39 $\rightarrow$ 30.55 2.81 $\rightarrow$ 3.02 53.22 $\rightarrow$ 60.47
Full Fine-tuning 11.93 $\rightarrow$ 12.24 30.86 $\rightarrow$ 31.38 3.21 $\rightarrow$ 3.25 63.38 $\rightarrow$ 65.21
LLaMA-3.1OCR Zero-shot Learning 4.37 $\rightarrow$ 4.59 21.42 $\rightarrow$ 21.89 1.50 $\rightarrow$ 1.72 34.06 $\rightarrow$ 40.24
QLoRA Fine-tuning 12.11 $\rightarrow$ 12.33 30.52 $\rightarrow$ 30.78 2.77 $\rightarrow$ 2.98 53.19 $\rightarrow$ 60.38
Full Fine-tuning 12.42 $\rightarrow$ 12.75 31.72 $\rightarrow$ 32.19 3.32 $\rightarrow$ 3.38 65.84 $\rightarrow$ 67.53
Qwen2-Audio Zero-shot Learning 4.29 $\rightarrow$ 4.51 21.53 $\rightarrow$ 22.18 1.59 $\rightarrow$ 1.77 34.31 $\rightarrow$ 40.52
QLoRA Fine-tuning 12.05 $\rightarrow$ 12.19 30.77 $\rightarrow$ 31.04 2.80 $\rightarrow$ 3.01 54.27 $\rightarrow$ 61.44
Full Fine-tuning 12.37 $\rightarrow$ 12.68 31.63 $\rightarrow$ 32.12 3.22 $\rightarrow$ 3.25 66.62 $\rightarrow$ 68.25
Table 6: Performance of baseline vs. planning models in non-video settings across different learning regimes. Each cell shows the result before $\rightarrow$ after applying the planning method.
Notably, despite these gains, video-based planning models such as Plan-mPLUG-Owl3 still outperform their non-visual counterparts by a notable margin. Nonetheless, our findings reinforce the idea that structured planning improves summarization quality beyond the video domain. In Appendix H, we further explore the effect of video content on our summarization task, varying the length of the video given as input to the model. We also perform experiments with different textual contexts for generating plan questions in Appendix I, and with controlled generation in Appendix J. Additionally, we present an error analysis of model output in Appendix K.
@@ 7 Human Evaluation
We conduct a human evaluation on 50 randomly selected instances from the VISTA test set. Annotators include master’s and doctoral students in computer science or computational linguistics with advanced English proficiency. They receive compensation per our university’s standard rate and are blind to the source of each summary to ensure impartial assessment. We compare Plan-mPlug-Owl3, mPLUG-Owl3, LLAVA-NeXT-Interleave, and GPT-o1 against human reference summaries. Three independent annotators are asked to review the source video and evaluate corresponding model outputs (and the human upper bound) on a 1–5 Likert scale for Faithfulness, Relevance, Informativeness, Conciseness, and Coherence (higher scores indicate better quality). They are also asked to provide an overall ranking. In total, participants rated 750 samples ($50\times 5\times 3$). Appendix N contains the full evaluation instructions.
@@
Figure 6 presents the performance of each model, along with the proportion of instances where models are rated best or worst. Fleiss’ Kappa scores for Faithfulness ($\kappa=0.767$), Relevance ($\kappa=0.842$), Informativeness ($\kappa=0.721$), Conciseness ($\kappa=0.792$), and Coherence ($\kappa=0.813$) indicate a substantial level of agreement, with an average agreement score of $\kappa=0.787$. Overall, human-written summaries outperform all neural summarization models in quality, as they are perceived as substantially more faithful, coherent, concise, and informative. Human-written summaries are 81.7% more likely to be rated as best compared to model-generated summaries.
Among the four neural models, GPT-o1 performs worst, being rated as worst 63.2% of the time. LLAVA-NeXT-Interleave follows suit, with a 17.8% chance of receiving the worst ranking. The plan-based model, Plan-mPLUG-Owl3, outperforms mPLUG-Owl3 and demonstrates superior performance across all metrics. Additionally, it stands out among neural summarization systems for its higher likelihood of generating high-quality summaries. Paired t-tests show that human answers are considered significantly better than all neural models in all metrics ($p<0.05$), revealing a clear gap between automatic systems and human performance on the VISTA dataset. The plan-based method is significantly better ($p<0.05$) than other neural models in faithfulness, coherence, and informativeness, although it falls short of human performance. We also evaluate all samples of the test set with an LMM-as-Judge and obtain results that are broadly consistent with human evaluation. We describe the details of this study in Appendix L.
@@ 8 Conclusion
This paper introduces VISTA, a novel dataset specifically curated for the task of summarizing scientific video presentations into concise and coherent textual summaries. Comprehensive evaluations across multiple large (language/audio/multimodal) models demonstrate that this task poses significant challenges due to the complexity and multimodal nature of scientific presentations. To address these challenges, we operate a plan-based summarization approach that incorporates discourse-aware planning prior to summary generation. This method consistently improves summary quality, factual coverage, and coherence across multiple settings. In addition to presenting the dataset, our study reveals that even the strongest current models still fall short of matching human performance by a noticeable margin. We believe that VISTA could provide a robust and extensible foundation for future research on video-to-text summarization.
@@
All the summary and video data used in this study are open source. While our sources are generally of high quality and exhibit a broad range of diversity, we have not investigated inherent biases in the data. Moreover, as these data represent only a small fraction of real-world data, our findings may not extend to all video-to-text summarization scenarios.
@@
In our task, we consider the paper abstract as a proxy for the summary of the corresponding video. This hypothesis has been supported by our two-stage quality control process, which ensures a strong alignment. However, we acknowledge that there may be nuanced differences between the abstract and a textual summary derived solely from the video. That said, authors often present the abstract as a summary of the video, as it conveys the key contributions, objectives, and findings of the research, which are typically central to the content discussed.
@@
We have tested the plan-based approach on the video-based, audio-based, and text-based large models in our experiments. Our work does not aim to prove that the plan-based method is effective in all models of different modalities. Moreover, plan-based methods can take many different forms, and our work does not aim to identify the optimal planning approach for our dataset.
@@
LMM-as-Judge
Although the LMM-based judge paradigm enables large-scale and relatively consistent evaluations, it may inherit biases from its pretraining data, and its black-box nature makes the rating process difficult to interpret. Data contamination also remains a concern if GPT-o1 is trained on overlapping data. We validate GPT-o1’s ratings with human evaluations on a small subset of samples, but this may not fully capture the model’s reliability across diverse topics, domains, or summary styles. Therefore, results should be interpreted with caution and supplemented by human judgment where possible.